# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v4

@@ 3.4. Per-Vector Dense-and-Sparse Quantization
As shown in Figure 5 in Appendix D, for both Keys and Values, the majority of elements are contained within a small percentage of the dynamic range. This means that by leveraging dense-and-sparse quantization, as demonstrated in (kim2023squeezellm, ), in order to isolate a small percentage of numerical outliers, we can restrict the range that we need to represent, thereby allowing us to represent the remaining elements with greater precision. However, when looking at the Key and Value distributions in Figure 2, different channels and tokens have different average magnitudes. Therefore, an element which counts as an outlier in one channel may not be an outlier in another channel (since that channel may have a greater average magnitude), making naive application of dense-and-sparse quantization suboptimal. It is therefore crucial to directly target the outlier values that skew the dynamic range at the granularity that we are quantizing in order to address the values that are exaggerating the range along that particular dimension. In this work, we leverage per-vector dense-and-sparse quantization, where we use a different outlier threshold per-vector (either a separate threshold per-channel for per-channel quantization, or a separate threshold per-token for per-token quantization), rather than a single outlier threshold for each layer.
@@
Note that computing outlier thresholds for per-vector dense-and-sparse quantization poses potential accuracy and efficiency challenges. However, in Section 3.6, we show that we are able to accurately calibrate for per-channel outlier thresholds offline and efficiently compute per-token outlier thresholds online. After determining the upper and lower outlier thresholds, the remaining numbers in the vector are normalized to the range $[-1,1]$, and we then minimize Equation 2 (ignoring outliers) in order to obtain the quantization signposts for the non-uniform datatype for the remaining numbers. Appendix H will demonstrate the benefits of removing a small percentage of numerical outliers and keeping them in full precision, as well as the advantages of per-vector dense-and-sparse quantization over using a single global outlier threshold for each layer. As shown in Figure 1, by removing 1% of numerical outliers using per-vector outlier thresholds, we achieve an additional 0.19 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization, which is within 0.07 perplexity of the fp16 baseline.
@@
Table 2. Evaluation of our method for different models using the perplexity (PPL) on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. KV cache sizes are estimated assuming a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention Sink-Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix N. Table 16 in Appendix M demonstrates a full evaluation on all LLaMA, Llama-2, Llama-3, and Mistral models.
@@
3.5. Attention Sink-Aware Quantization
Prior work has demonstrated that after the first few layers in LLMs, the model tends to allocate a large attention score to the first token (xiao2023efficient, ). This occurs even when the initial token is not semantically important. This phenomenon happens because the model tends to use the inital token as a “sink”. In our work, we demonstrate that due to the Attention Sink phenomenon, the model is disproportionately sensitive to quantization error in the first token. By keeping only the first token in fp16, we can attain perplexity benefits, particularly for 2-bit quantization. A similar observation has also been made in another concurrent work  (liu2024intactkv, ). Note that when retaining the first token in fp16, we account for this during the calibration process as well, meaning that we ignore the first token when deriving the nuqX datatype and when calibrating the scaling factors and zero points offline for the Keys. As demonstrated in Appendix I, this approach persistently yields performance benefits, particularly with lower bit widths and without dense-and-sparse quantization.
3.6. Offline Calibration versus Online Computation
A crucial challenge for activation quantization is that we either need to compute statistics on-the-fly (which is potentially expensive) or else we need to use offline calibration data (which potentially has negative accuracy implications). The challenges with computing scaling factors (and zero-points) online versus offline for both Keys and Values are shown in Figure 3. In per-channel quantization, it is challenging to update scaling factors online since the scaling factors corresponding to each incoming channel would potentially need to be updated whenever a new token is added to the KV cache. It is therefore desirable to be able to compute statistics offline (i.e., using calibration data before running inference). While this can have negative effects on model accuracy, in Appendix J we show that we can effectively calibrate offline for per-channel quantization, obviating the need for online updates of scaling factors for per-channel quantization. For per-token quantization, it is challenging to calibrate for scaling factors offline due to the presence of outlier Value tokens. It is therefore desirable to be able to compute scaling factors and outlier thresholds online for each incoming token. As shown in Appendix J, we can efficiently compute outlier thresholds online per-token by offloading to the CPU. By leveraging custom quantization function implementations for compressing activations, we are able to perform online per-token Value quantization without compromising on performance.
3.7. Kernel Implementation
In order to efficiently perform activation quantization on-the-fly, we leverage dedicated kernel implementations with our 4-bit quantization method for compressing vectors to reduced precision and extracting the sparse outliers, performing matrix-vector multiplications using the compressed vectors, and performing sparse matrix-dense vector multiplications using the sparse outliers. We store the quantized Key and Value matrices as 4-bit elements which are used as indices into lookup tables to recover the corresponding fp16 values. We store the sparse outlier matrices in either Compressed-Sparse Row (CSR) or Compressed-Sparse Column (CSC) format (depending on which aligns better with appending new Key and Value tokens). The kernels for the Key matrix-vector operations apply RoPE on-the-fly in order to support pre-RoPE quantization. More kernel implementation details are provided in Appendix P.
Figure 4. Perplexity results for the LLaMA-2-7B-32K model (chen2023extending, ) as well as the Llama-2-70B-32K LongLoRA model (chen2023longlora, ) on the Wikitext-2 dataset, evaluated using different sequence lengths.
4. Results
4.1. Main Evaluation
@@
We used the LLaMA-7B/13B/30B/65B, Llama-2-7B/13B/70B, Llama-3-8B/70B, and Mistral-7B models to evaluate our methodology by measuring perplexity on both Wikitext-2 and C4 (touvron2023llama, ; touvron2023llama2, ; llama3modelcard, ; jiang2023mistral, ; wikitext2, ; c4, ); see Appendix K for details on our experimental setup, as well as our methodology for computing KV cache size estimates. Table 2 shows the results for LLaMA models for the Wikitext-2 dataset. We compared our method with per-token quantization with and without grouping. The baseline configurations used by Atom and FlexGen are included for reference (zhao2023atom, ; sheng2023flexgen, ). We find that our method consistently outperforms baseline approaches by an especially large margin with 3-bit and 2-bit quantization. Once we incorporate outliers, we further push the performance of low-precision quantization, achieving 4-bit quantization with less than 0.02 perplexity degradation, 3-bit quantization with under 0.1 perplexity degradation, and 2-bit quantization with under 0.5 perplexity degradation on Wikitext-2, relative to the fp16 baseline, across all models (while attaining 3.7$\times$, 4.8$\times$, and 6.9$\times$ memory savings, respectively).
4.2. Long Context Length Evaluation
We evaluated long context length performance using the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation (chen2023extending, )) as well as the Llama-2-70B-32K LongLoRA model (chen2023longlora, ). For evaluating performance on longer context lengths, we first evaluated perplexity on Wikitext-2 using larger amounts of input context, as shown in Figure 4 (chen2023longlora, ; han2023lminfinite, ). The results demonstrate how our method maintains accuracy even for longer amounts of input context, thereby enabling efficient and accurate long sequence length inference.
We also evaluated the performance of our quantization method on passkey retrieval to assess the model’s ability to use its context. Passkey retrieval involves evaluating the model’s capacity to locate specific information in long texts (longchat2023, ), and this can be used to effectively measure the maximum distance over which a token can attend during the inference stage. We used the passkey evaluation framework from (zhu2023pose, ) (which is based on the methodology from (mohtashami2023landmark, )) to evaluate retrieval performance. The passkey retrieval results are provided in Table 3, demonstrating how our method is able to maintain the retrieval performance for long context length models. We also include comparisons with the passkey retrieval when using KIVI for the LLaMA-2-7B-32K model (kivi, ), demonstrating that our approach can attain higher retrieval rate for the same compression level. Our improved performance on retrieval tasks can be attributed to our improved representation of all tokens equally (whereas KIVI preserves a local window of residual tokens in fp16, which better represents the tail part of the context but doesn’t provide as much benefit for tasks that need to utilize the full context window). Additionally, we include evaluation on Longbench in Appendix O, which contains a suite of long-context length evaluation benchmarks including QA tasks, summarization, and few-shot learning (bai2023longbench, ). We demonstrate that our 3-bit approach can attain minimal degradation relative to fp16 inference, and that our approach outperforms KIVI for a similar compression level. Note that for generative tasks, when processing the input prompt, the Key/Value matrix multiplications are computed using the fp16 Keys and Values, and then the Keys and Values are separately compressed into low precision.
@@ 4.2. Long Context Length Evaluation
Table 3. Passkey retrieval results across different context lengths for the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation (chen2023extending, )) as well as the Llama-2-70B-32K LongLoRA model (chen2023longlora, ). The values reported are the success rate for retrieving the passkey, computed over 50 samples. We also include comparisons with KIVI for reference, using the 2-bit configuration with group size of 32 and 128-element fp16 residual (kivi, ). Average bit widths are estimated for each approach assuming 32K context length. Note that the open-source code for running KIVI with LLaMA does not support grouped-query attention, so we did not include comparisons with KIVI for Llama-2-70B-32K.
Model Method 2K 4K 8K 16K 32K Bit Width
@@
4.3. Joint Weight and KV Cache Quantization
Table 4 provides results for our method when applied along with quantizing weights to low precision (where the weight quantization follows the approach in (kim2023squeezellm, )). We observe minimal perplexity degradation when leveraging our KV cache quantization approach, even when weights are also quantized to reduced precision (achieving within 0.02 and 0.1 perplexity of 4-bit and 3-bit weight-only quantization, respectively, when quantizing the KV cache using nuq4-1% for the LLaMA-7B and LLaMA-13B models). These results demonstrate how our method is compatible with existing weight-only quantization methods.
Table 4. KV cache quantization results when KVQuant is applied in conjunction with the weight quantization methodology in SqueezeLLM (kim2023squeezellm, ). w4-s45 and w3-s45 for weights refer to the 4-bit and 3-bit dense-and-sparse weight quantization approaches in  (kim2023squeezellm, ), respectively. See Appendix K for experimental details.
@@ 4.3. Joint Weight and KV Cache Quantization
4.4. Performance Analysis and Memory Savings
@@
Table 5. Average latency (in microseconds) for the Key and Value nuq4-1% kernels, benchmarked on an A6000 GPU for the LLaMA-2-7B-32K model across different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the fp16 Key multiplication time also includes the time to apply RoPE to the newly appended Key vector.  Section 3.7 and Appendix P provide additional details for our kernel implementation, Appendix P describes our benchmarking methodology, and Table 20 provides a detailed breakdown of kernel runtime on an A6000 GPU.
@@
Table 6. Activation memory size estimates (GB) for 128K, 1M, and 10M sequence length ($l$) for different LLaMA models. By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.

### v5

@@ 3.4 Per-Vector Dense-and-Sparse Quantization
As shown in Figure 4 in Appendix F, for both Keys and Values, the majority of elements are contained within a small percentage of the dynamic range. This means that by leveraging dense-and-sparse quantization, as demonstrated in [17], in order to isolate a small percentage of numerical outliers, we can restrict the range that we need to represent, thereby allowing us to represent the remaining elements with greater precision. However, when looking at the Key and Value distributions in Figure 2, different channels and tokens have different average magnitudes. Therefore, an element which counts as an outlier in one channel may not be an outlier in another channel (since that channel may have a greater average magnitude), making naive application of dense-and-sparse quantization suboptimal. It is therefore crucial to directly target the outlier values that skew the dynamic range at the granularity that we are quantizing in order to address the values that are exaggerating the range along that particular dimension. In this work, we leverage per-vector dense-and-sparse quantization, where we use a different outlier threshold per-vector (either a separate threshold per-channel for per-channel quantization, or a separate threshold per-token for per-token quantization), rather than a single outlier threshold for each layer.
@@
Note that computing outlier thresholds for per-vector dense-and-sparse quantization poses potential accuracy and efficiency challenges. However, in Section 3.6, we show that we are able to accurately calibrate for per-channel outlier thresholds offline and efficiently compute per-token outlier thresholds online. After determining the upper and lower outlier thresholds, the remaining numbers in the vector are normalized to the range $[-1,1]$, and we then minimize Equation 1 (ignoring outliers) in order to obtain the quantization signposts for the non-uniform datatype for the remaining numbers. Appendix J will demonstrate the benefits of removing a small percentage of numerical outliers and keeping them in full precision, as well as the advantages of per-vector dense-and-sparse quantization over using a single global outlier threshold for each layer. As shown in Figure 1, by removing 1% of numerical outliers using per-vector outlier thresholds, we achieve an additional 0.19 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization, which is within 0.07 perplexity of the fp16 baseline.
3.5 Attention Sink-Aware Quantization
Prior work has demonstrated that after the first few layers in LLMs, the model tends to allocate a large attention score to the first token [42]. This occurs even when the initial token is not semantically important. This phenomenon happens because the model tends to use the inital token as a “sink”. In our work, we demonstrate that due to the Attention Sink phenomenon, the model is disproportionately sensitive to quantization error in the first token. By keeping only the first token in fp16, we can attain perplexity benefits, particularly for 2-bit quantization. A similar observation has also been made in another concurrent work  [23]. Note that when retaining the first token in fp16, we account for this during the calibration process as well, meaning that we ignore the first token when deriving the nuqX datatype and when calibrating the scaling factors and zero points offline for the Keys. As demonstrated in Appendix K, this approach persistently yields performance benefits, particularly with lower bit widths and without dense-and-sparse quantization.
3.6 Offline Calibration versus Online Computation
A crucial challenge for activation quantization is that we either need to compute statistics on-the-fly (which is potentially expensive) or else we need to use offline calibration data (which potentially has negative accuracy implications). The challenges with computing scaling factors (and zero-points) online versus offline for both Keys and Values are shown in Figure 5 in Appendix L. In per-channel quantization, it is challenging to update scaling factors online since the scaling factors corresponding to each incoming channel would potentially need to be updated whenever a new token is added to the KV cache. It is therefore desirable to be able to compute statistics offline (i.e., using calibration data before running inference). While this can have negative effects on model accuracy, in Appendix L we show that we can effectively calibrate offline for per-channel quantization, obviating the need for online updates of scaling factors for per-channel quantization. For per-token quantization, it is challenging to calibrate for scaling factors offline due to the presence of outlier Value tokens. It is therefore desirable to be able to compute scaling factors and outlier thresholds online for each incoming token. As shown in Appendix L, we can efficiently compute outlier thresholds online per-token by offloading to the CPU. By leveraging custom quantization function implementations for compressing activations, we are able to perform online per-token Value quantization without compromising on performance.
3.7 Kernel Implementation
In order to efficiently perform activation quantization on-the-fly, we leverage dedicated kernel implementations with our 4-bit quantization method for compressing vectors to reduced precision and extracting the sparse outliers, performing matrix-vector multiplications using the compressed vectors, and performing sparse matrix-dense vector multiplications using the sparse outliers. We store the quantized Key and Value matrices as 4-bit elements which are used as indices into lookup tables to recover the corresponding fp16 values. We store the sparse outlier matrices in either Compressed-Sparse Row (CSR) or Compressed-Sparse Column (CSC) format (depending on which aligns better with appending new Key and Value tokens). The kernels for the Key matrix-vector operations apply RoPE on-the-fly in order to support pre-RoPE quantization. More kernel implementation details are provided in Appendix R.
@@ 3.7 Kernel Implementation
Table 1: Evaluation of our method for different models using the perplexity (PPL) on Wikitext-2. KVQuant results are using pre-RoPE per-channel quantization for Keys. KV cache sizes are estimated assuming a sequence length of 128K (ignoring context length limits for the models). Note that ATOM and FlexGen use 4-bit quantization with group sizes of 128 and 64 with uniform quantization, respectively, and we extend their methods to 3-bit and 2-bit quantization. We leverage Attention Sink-Aware quantization for all bit widths. We used post-RoPE quantization for all baseline methods since it achieves higher accuracy when quantizing Keys per-token as shown in Appendix P. Table 18 in Appendix O demonstrates a full evaluation on all LLaMA, Llama-2, Llama-3, and Mistral models.
@@
4 Results
4.1 Main Evaluation
We used the LLaMA-7B/13B/30B/65B, Llama-2-7B/13B/70B, Llama-3-8B/70B, and Mistral-7B models to evaluate our methodology by measuring perplexity on both Wikitext-2 and C4 [36, 37, 1, 16, 27, 31]. Perplexity has been measured using teacher forcing with the output logits of all input tokens. We compared our method against (i) uniform quantization without grouping (intX), (ii) nonuniform quantization using NormalFloat [8] without grouping (nfX), as well as (iii) Atom [44] and FlexGen [34]. Note that Atom and FlexGen use uniform quantization with group sizes of 64 and 128, respectively. All the KVQuant models throughout this experiment section are calibrated using 16 calibration samples of sequence length 2K from the Wikitext-2 training set. See Appendix M for details on our experimental setup, including our methodology for computing KV cache size estimates.
@@ 4.1 Main Evaluation
Table 1 shows the results for LLaMA models for the Wikitext-2 dataset. We compared our method with per-token quantization with and without grouping. The baseline configurations used by Atom and FlexGen are included for reference [44, 34]. We find that our method consistently outperforms baseline approaches by an especially large margin with 3-bit and 2-bit quantization. Once we incorporate outliers, we further push the performance of low-precision quantization, achieving 4-bit quantization with less than 0.02 perplexity degradation, 3-bit quantization with under 0.1 perplexity degradation, and 2-bit quantization with under 0.5 perplexity degradation on Wikitext-2, relative to the fp16 baseline, across all models (while attaining 3.7$\times$, 4.8$\times$, and 6.9$\times$ memory savings, respectively).
Figure 3: Perplexity results for the LLaMA-2-7B-32K model [5] as well as the Llama-2-70B-32K LongLoRA model [6] on the Wikitext-2 dataset, evaluated using different sequence lengths.
@@
Table 2: Passkey retrieval results across different context lengths for the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation [5]) as well as the Llama-2-70B-32K LongLoRA model [6]. The values reported are the success rate for retrieving the passkey, computed over 50 samples. We also include comparisons with KIVI for reference, using the 2-bit configuration with group size of 32 and 128-element fp16 residual [26]. Average bit widths are estimated for each approach assuming 32K context length. Note that the open-source code for running KIVI with LLaMA does not support grouped-query attention, so we did not include comparisons with KIVI for Llama-2-70B-32K.
Model Method 2K 4K 8K 16K 32K Avg. Bit Width
@@
Table 3: LongBench evaluation for the LLaMA-2-7B-32K model using KVQuant-3bit-1%. Comparisons with KIVI are included for reference, using the configuration with group size of 32 and 128-element fp16 residual [26]. Average bit widths are estimated for each approach assuming 12.2K context length, which was the average number of tokens across all tasks.
Config
Avg. bit
NtrvQA
Qasper MF-en Hotpot 2Wiki Musique GovRep QMSum MNews TREC TriviQA SamSum RBench LCC PsgRetr PsgCnt Avg.
fp16 Baseline 16 17.96 10.51 33.43 12.55 12.53 6.19 29.65 16.99 22.15 71 87.79 43.97 59.99 62.14 23 1.50 31.96
KIVI-2-gs32-r128 3.17 19.25 10.66 24.78 12.48 11.19 6.38 27.05 16.36 23.37 71 80.80 43.93 57.74 60.61 13.58 1.50 30.04
KVQuant-3bit-1% 3.33 18.87 13.67 30.93 12.07 12.55 6.25 27.10 16.53 16.54 71 87.55 43.95 59.50 61.52 19.5 1.75 31.21
Table 4: RULER evaluation results for the LLaMA-2-7B-32K model with KVQuant quantization methods. We report accuracy across RULER tasks, comparing our KVQuant configurations to baseline and KIVI approaches. A maximum context length of 32K is used for evaluation. Our results show that our method retains baseline accuracy even with aggressive quantization and pruning.
Config
Avg. bit
Niah1
Niah2 Niah3 MKey1 MKey2 MKey3 MValue MQuery VT CWE FWE QA1 QA2 Avg.
fp16 Baseline 16 100 99.8 98.6 94 68.2 11 55.95 64.5 37.88 9.64 30.4 31.6 31.6 56.40
KIVI-2-gs32-r128 3.05 76 85.6 59.6 72.6 11.4 0 34.7 46.45 39.6 8.26 30.53 24.8 27.6 39.78
KVQuant-3bit-1% 3.33 99.8 98.8 95.2 92.8 61.6 6.4 47.5 54.45 41.04 8.52 29.33 31.0 31.0 53.65
KVQuant-2bit-1% 2.33 95.4 86.8 49.8 73.6 23.4 0 16.65 22.95 22.52 5.14 24.0 26.4 28.4 36.54
Table 5: KV cache quantization results when KVQuant is applied in conjunction with the weight quantization methodology in SqueezeLLM [17]. w4-s45 and w3-s45 for weights refer to the 4-bit and 3-bit dense-and-sparse weight quantization approaches in  [17], respectively. See Appendix M for experimental details.
@@
4.2 Long Context Length Evaluation
Perplexity Evaluation. We evaluated long context length performance using the LLaMA-2-7B-32K model (uptrained for long sequence lengths using positional interpolation [5]) as well as the Llama-2-70B-32K LongLoRA model [6]. For evaluating performance on longer context lengths, we first evaluated perplexity on Wikitext-2 using larger amounts of input context, as shown in Figure 3 [6, 13]. The results demonstrate how our method maintains accuracy even for longer amounts of input context, thereby enabling efficient and accurate long sequence length inference.
Passkey Retrieval Evaluation. We also evaluated the performance of our quantization method on passkey retrieval to assess the model’s ability to use its context. Passkey retrieval involves evaluating the model’s capacity to locate specific information in long texts [19], and this can be used to effectively measure the maximum distance over which a token can attend during the inference stage. We used the passkey evaluation framework from [45] (which is based on the methodology from [28]) to evaluate retrieval performance. The passkey retrieval results are provided in Table 2, demonstrating how our method is able to maintain the retrieval performance for long context length models. We also include comparisons with the passkey retrieval when using KIVI for the LLaMA-2-7B-32K model [26], demonstrating that our approach can attain higher retrieval rate for the same compression level. Our improved performance on retrieval tasks can be attributed to our improved representation of all tokens equally. This approach differs from KIVI, which preserves a local window of residual tokens in fp16. Therefore, while KIVI is effective at representing the tail part of the context, it may provide less benefit for tasks requiring the utilization of the full context window.
LongBench Evaluation. Table 3 shows evaluation on LongBench [3] for the LLaMA-2-7B-32K model. LongBench contains a suite of long-context length evaluation benchmarks including QA tasks, summarization, and few-shot learning [3]. The max input context length is set at 31500, and results using KIVI are also included for reference [26]. Our results demonstrate that our 3-bit model can attain minimal degradation relative to the fp16 baseline, outperforming KIVI for a similar compression level.
RULER Evaluation. Finally in Table 4, we provide evaluation of KVQuant and KIVI on the RULER benchmark suite [15] using LLaMA-2-7B-32K. As can be seen in the table, 3-bit KVQuant achieves 14% better score against KIVI with a similar average bit-width. Furthermore, our 2-bit KVQuant achieves similar accuracy to KIVI with 1.5$\times$ smaller bit-width.
4.3 Joint Weight and KV Cache Quantization
Table 5 provides results for our KV cache quantization method when the weights are also quantized using the methodology in SqueezeLLM [17]. We observe minimal perplexity degradation when leveraging our KV cache quantization approach, even when weights are also quantized to reduced precision. In particular, we observe small 0.02 and 0.1 perplexity degradation of 4-bit and 3-bit weight-only quantization, respectively, when quantizing the KV cache using nuq4-1% for the LLaMA-7B and LLaMA-13B models. These results demonstrate how our method is compatible with existing weight-only quantization methods.
4.4 Performance Analysis and Memory Savings
@@
Appendix A highlights the benefits of KVQuant in supporting longer context lengths through reducing KV cache memory footprint. As shown in Table 8 in Appendix A, our nuq2 method provides 8$\times$ KV cache compression and enables serving the quantized LLaMA-7B model with a context length of 1M tokens on a single A100 GPU, as well as enabling serving the LLaMA-7B model with 10M context length on an 8-GPU system. Our results show little degradation compared to baseline fp16 inference while providing significant compression, demonstrating the benefits of our approach for enabling accurate and efficient long sequence length inference.
Table 6: Average latency (in microseconds) for the Key and Value nuq4-1% kernels, benchmarked on an A6000 GPU for the LLaMA-2-7B-32K model across different sequence lengths ($l$). fp16 matrix-vector multiplication latencies are included for reference, and the fp16 Key multiplication time also includes the time to apply RoPE to the newly appended Key vector.  Section 3.7 and Appendix R provide additional details for our kernel implementation, Appendix R describes our benchmarking methodology, and Table 22 provides a detailed breakdown of kernel runtime on an A6000 GPU.
@@
5 Conclusion
Acknowledgements
The authors would like to acknowledge Nicholas Lee for helpful discussions and feedback. We acknowledge gracious support from Intel, Furiosa, Apple, Samsung SAIT, and NVIDIA. We also appreciate the support from Microsoft through their Accelerating Foundation Model Research, including great support from Sean Kuno. Furthermore, we appreciate support from Google Cloud, the Google TRC team, and specifically Jonathan Caton, and Prof. David Patterson. Prof. Keutzer’s lab is sponsored by the Intel corporation, Intel One-API, Intel VLAB team, the Intel One-API center of excellence, as well as funding through BDD and BAIR. We appreciate great feedback and support from Ellick Chan, Saurabh Tangri, Andres Rodriguez, and Kittur Ganesh. Sehoon Kim would like to acknowledge the support from the Korea Foundation for Advanced Studies (KFAS). Michael W. Mahoney would also like to acknowledge a J. P. Morgan Chase Faculty Research Award as well as the DOE, NSF, and ONR. Our conclusions do not necessarily reflect the position or the policy of our sponsors, and no official endorsement should be inferred.
Appendix A Memory Bottlenecks for Long Context Length Inference
Table 7 shows the model size and KV cache memory requirements for different LLaMA models with different sequence lengths. For short sequence lengths, the model weights are the primary memory bottleneck. However, for longer sequence lengths and larger batch sizes, the KV cache memory is the main bottleneck. This is particularly pronounced when the weights are already quantized to low precision. In our work, we demonstrate that we can help address the KV cache memory bottleneck through low-precision KV cache quantization. By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.
Table 7: Model size and activation memory size estimates for different sequence lengths and batch sizes (BS) for different LLaMA models. For long sequence lengths and larger batch sizes, activation memory is the main bottleneck (particularly when weights are already quantized to low precision). By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.
BS Model Model Size (GB) KV Cache Size w/ Diff. Seq Len (GB)
16 $\rightarrow$ 2-bit 32K 128K 1M 10M (16 $\rightarrow$ 2-bit)
1 7B 12.6 $\rightarrow$ 1.6 16 64 512 4883 $\rightarrow$ 610
13B 24.1 $\rightarrow$ 3.0 25 100 800 7629 $\rightarrow$ 954
30B 60.3 $\rightarrow$ 7.5 49 195 1560 14877 $\rightarrow$ 1860
65B 121.1 $\rightarrow$ 15.1 80 320 2560 24414 $\rightarrow$ 3052
4 7B 12.6 $\rightarrow$ 1.6 64 256 2048 19531 $\rightarrow$ 2441
13B 24.1 $\rightarrow$ 3.0 100 400 3200 30518 $\rightarrow$ 3815
30B 60.3 $\rightarrow$ 7.5 195 780 6240 59509 $\rightarrow$ 7439
65B 121.1 $\rightarrow$ 15.1 320 1280 10240 97656 $\rightarrow$ 12207
Table 8: Activation memory size estimates (GB) for 128K, 1M, and 10M sequence length ($l$) for different LLaMA models. By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.